Dan is gliding on his skateboard at 2.00 m/s . He suddenly jumps backward off the skateboard, kicking the skateboard forward at 6.00 m/s . Dan's mass is 70.0 kg and the skateboard's mass is 6.00 kg . How fast is Dan going as his feet hit the ground?
Answer #1

Using conservation momentum

MV=m1v1+m2v2 .....1

here M=70+6=76kg, m1=70kg, m2=6kg ,v2=6m/s

put these values in equation ..1

76*2=70v1 +6*6

v1=(152-36)/70

=1.66 m/s answer